OpenStudy (shaw74):

An angle measure of 82 degrees is equivalent to ____ radians. Round your answer to the nearest hundredth when necessary.

OpenStudy (shaw74):

im lost

OpenStudy (shaw74):

@study100

OpenStudy (shaw74):

@SolomonZelman

OpenStudy (shaw74):

@johnweldon1993

OpenStudy (shaw74):

@anyboddy

OpenStudy (johnweldon1993):

To convert to radians from degrees you multiply by pi/180 for example \[\large 30^\circ \times \frac{\pi}{180^\circ} = \frac{30\pi}{180} = \frac{\pi}{6}\text{radians}\]

OpenStudy (cwrw238):

use the equivalence 180 degrees = pi radians so 1 degree = pi/180 radians therefore 82 degrees = ? radians

OpenStudy (shaw74):

6 radians

OpenStudy (cwrw238):